The mean salary offered to students who are graduating from Coastal State University this year is $24,245 , with a standard deviation of $3643. A random sample of 80 Coastal State students graduating this year has been selected. What is the probability that the mean salary offer for these 80 students is $24,250 or less?
1
Expert's answer
2020-07-08T18:45:48-0400

Given : mean= 24245, s= 3643,n= 80


To find : P(X"\leq" 24250)


solution:

P("\overline{X}" ≤ 24250)= P("\frac{\overline{X}-mean}{\frac{s}{\sqrt{n}}}\leq\frac{24250-24245}{\frac{3643}{\sqrt{80}}} )"

= P(Z"\leq" 0.0123)


Using Z table ,"P(Z\leq 0.0123)" = 0.5049.


Answer: The probability that the mean salary offer for these 80 students is $24,250 or less will be 0.5049.